# The Convergence Frontier: Integrating Machine Learning and High Performance Quantum Computing for Next-Generation Drug Discovery

### 18 Mar 2026 v1

@@ 1 Introduction
This here paper is organized to follow a streamlined, end-to-end drug design workflow utilizing both classical and quantum computing platforms. We begin by discussing the preparation of molecular systems, including the water placement problem in quantum computing. We then analyze quantum algorithms for quantum chemistry as providers of high-accuracy training data for quantum-based Neural Network Potentials. A focus on the Hyperion quantum emulator is provided alongside the discussion on quantum hardware. We then focus on the development of these specific Machine Learning models for Molecular Dynamics simulations, introducing the FeNNix-Bio1 foundation model and its underlying Ignis database, and showing the transition from classical to exact quantum algorithms for training data generation. Finally, we provide an overview of modern enhanced sampling techniques and their projected quantum speedups. Throughout each section, we contrast classical and quantum implementations while providing detailed performance and resource estimations.
@@ 3.2 Quantum Computing: Overview of Recent Quantum Algorithms for Electronic Structure
3.2.1 Near-Term Hybrid Quantum Algorithms: the Search for Hardware Efficient Ansätze
The VQE approximates the molecular ground state by minimizing the energy over a truncated subspace of the Full-CI Hilbert space, in close analogy with classical CI methods. A key difference is that VQE constructs this variational space using an ansatz composed of unitary, exponentially parameterized operators, whereas CI relies on linear expansions in excitation operators. Quantum advantage with VQE would then mean that, for some practically relevant problems, the variational space spanned by the quantum circuit provides a better heuristic than classical ansätze [65], assuming efficient parameter training. Then, the central challenge in VQE is ansatz design: it must balance expressivity (containing states near the ground state), trainability (enabling efficient parameter optimization), and circuit depth, as NISQ devices allow only limited numbers of qubits and gates due to decoherence. So far, practical implementations continue to face numerous challenges, several of which are outlined below. First, commonly used ansätze can be insufficiently expressive for strongly correlated systems, particularly when higher-order excitations such as triples are required to achieve chemical accuracy. Second, the operator pools from which ansätze are constructed are often far larger than algebraically necessary, resulting in substantial computational overhead. Last but not least, VQE relies on a global, non-convex optimization landscape that is often rough and largely flat, due both to barren plateaus arising from high dimensionality and to excited-state contamination intrinsic to energy minimization, leading to significant convergence difficulties in multi-parameter optimization.
@@ 3.2.1 Near-Term Hybrid Quantum Algorithms: the Search for Hardware Efficient Ansätze
First, we focused on the fixed ansatz method utilizes a pre-defined quantum circuit structure, i.e. a template of gates, where the sequence and connectivity of the qubits remain constant throughout the computation. While the arrangement of the gates is static, the individual rotation angles (the parameters) are tunable. During the VQE process, a classical optimizer iteratively adjusts these angles to minimize the expectation value of the Hamiltonian, effectively searching for the system’s ground state energy within a constrained search space defined by that specific circuit.To improve expressivity in strongly correlated regimes, we extended the conventional chemistry-inspired Coupled Cluster UCCSD fixed ansatz [66] by incorporating triple excitations, leading to a UCCSDT formulation [67] capable of capturing higher-order correlation effects required for chemical accuracy. Second, to mitigate the inefficiency of oversized operator pools, the non-iterative disentangled UCC ansatz was introduced (NI-DUCC [68]) using a highly expressive network of hardware-efficient operators constructed to guarantee algebraic completeness. Third, we proposed an efficient and systematic procedure to generate complete operator pools associated with a given Lie algebra [69], which was previously a technically demanding task limiting qubit counts. This advance enables the practical deployment of minimal complete pools (MCPs) within NI-DUCC, i.e. NI-DUCC(MCPs).
Table 1: Comparison of Standard UCC-VQE and NI-DUCC(MCPs)-VQE
Feature Standard UCC-VQE NI-DUCC(MCPs)-VQE
Circuit Depth High (due to Trotterization) Low (Disentangled/Product form)
Optimization Gradient-based (Slow) Gradient-independent / Non-iterative
Symmetry Hard to maintain exactly Built-in via Lie Algebraic closure
Accuracy Limited to double excitations (no Full-CI accuracy) and by noise/depth on current hardware Reaches Full-CI at specific layers but still limited by noise/depth on current hardware
@@
In practice, MCPs can also be used within adaptive VQE frameworks such as ADAPT-VQE [70, 71]. Indeed, these latter have the ability to construct compact, problem-specific circuits that are significantly more efficient than fixed-structure hardware heuristics as they grow the circuit dynamically by adding operators one by one based on their impact. Consequently, MCPs underpin our introduction of MB-ADAPT-VQE technique [69], which couples MCPs with batch-ADAPT techniques[72] to ensure optimal convergence. This approach significantly enhances hardware efficiency by mitigating the high cost of gradients, a notorious bottleneck in adaptive methods.
Table 2: Comparison of ADAPT-VQE and MB-ADAPT-VQE
Feature Standard ADAPT-VQE MB-ADAPT-VQE
Operator Pool Typically uses large pools, such as the QEB pool. Uses optimally constructed Minimal Complete Pools (MCPs).
Pool Scaling Traditionally scales as $\mathcal{O}(N^{4})$. Verified to scale linearly, $\mathcal{O}(N)$, with the number of qubits.
Ansatz Growth Appends a single exponentiated generator per iteration. Incorporates MCPs into a batched approach, adding $k$ operators per iteration.
Selection Cost Requires measuring energy gradients for every operator in a large pool. Reduces measurement overhead by using a compact, symmetry-preserving pool
Gradient Cost Must evaluate energy gradients for $\mathcal{O}(N^{4})$ operators per step. Only requires gradient evaluations for $\mathcal{O}(N)$ operators per step.
Measurement Overhead High total measurement cost of approximately $\mathcal{O}(N^{8})$. Dramatically reduced measurement overhead scaling as $\mathcal{O}(N^{5})$.
Accuracy Reaches Full-CI at specific layers but still limited by noise/depth on current hardware Reaches Full-CI at specific layers but still limited by noise/depth on current hardware
By implementing NI-DUCC(MCPs) and MB-ADAPT-VQE, we demonstrate algorithmic parity with, or superiority over, current best-in-class fixed and adaptive VQE approaches for quantum chemistry.
Finally, to address excited-state contamination and the resulting energy plateaus in ADAPT-VQE, we introduced Overlap-ADAPT-VQE [73], which grows the ansatz via variational projection onto an intermediate-quality reference state rather than direct energy minimization. This projection-based strategy yields significantly more compact circuits and accelerates subsequent convergence. More broadly, Overlap-ADAPT-VQE establishes a unified framework for learning and compressing quantum states into efficient ansätze. We exploited this capability for classical data loading by compressing sparse selected-CI wavefunctions into compact quantum representations that serve as high-quality initial states for both VQE and Quantum Phase Estimation [73, 74], two algorithms whose performance is critically dependent on initialization quality.
Table 3: Comparison of Standard and Overlap-ADAPT-VQE Strategies
Feature Standard ADAPT-VQE Overlap-ADAPT-VQE
Growth Metric Energy Gradient (aims to minimize ground state energy). Overlap Maximization (aims to match a target wavefunction).
Target State The exact ground state of the system Hamiltonian. A high-quality intermediate target (e.g., from Selected-CI).
Ansatz Efficiency Can suffer from local minima and over-parameterization in noisy landscapes. Produces ”ultra-compact” ansätze by avoiding energy-landscape noise.
Note that while MB-ADAPT-VQE is highly efficient in terms of measurement overhead and convergence rate, it can be couple to Overlap-ADAPT to produce more compact representations.
@@
To conclude, let us discuss the practical difficulties of implementing VQE algorithms on todays quantum computers. To this end and in order to mitigate the optimization difficulties of VQE arising from non-convex landscapes and barren plateaus, we introduced Greedy Gradient-Free ADAPT-VQE (GGA-VQE) [75]. The method exploits analytical structure to convert the global variational optimization into an iterative greedy procedure in which each parameter can be determined locally without gradient estimation. This substantially reduces the classical optimization overhead while producing compact, hardware-aligned ansätze tailored to the target Hamiltonian.
Table 4: Comparison of Standard ADAPT-VQE and GGA-VQE Strategies
Feature Standard ADAPT-VQE GGA-VQE (Greedy Gradient-Free)
Selection Criterion Uses energy gradients to identify the most impactful operator. Uses a gradient-free greedy selection rule to pick operators.
Measurement Cost High; requires evaluating gradients for all pool operators at each step. Significantly lower; bypasses the hardware-intensive gradient-evaluation stage.
Optimization Iterative classical-quantum loop based on steepest descent. Aimed at reducing classical optimization overhead by avoiding gradient maps.
Scaling Issue Suffers from the ”gradient measurement bottleneck” as pools grow. Mitigates the bottleneck, though potentially at the cost of ansatz compactness.
Hardware Fit Limited by the ”gradient measurement bottleneck” on NISQ devices More hardware-efficient for intermediate-scale systems due to lower overhead.
@@
Despite significant fluctuations in the QPU-measured energies, the parameters extracted from hardware measurements remain close to the optimal noiseless values for the first iterations of the algorithm (Fig. 6), indicating that the greedy analytical updates are robust to realistic noise levels. At the same time, recent analyses by Dalton et al. [76] show that achieving a genuine quantum advantage with adaptive VQE would require several orders of magnitude reduction in current gate error rates. While GGA-VQE can be viewed as the ”minimal” adaptive method aimed at hardware implementation, primarily because it removes the most expensive part of the ADAPT-VQE process, i.e. the hardware-intensive gradient evaluation, MB-ADAPT-VQE and NI-DUCC-VQE offer more refined ”minimal” strategies by optimizing the underlying mathematical structure of the operator pools. In any case, taken together, these results reflect the current frontier of adaptive VQE on NISQ hardware, where algorithmic progress is possible but ultimately constrained by present noise levels. In practice, low-accuracy VQE computations can, of course, be performed on hardware with minimal basis sets. However, these do no meet yet the accuracy requirements to be included into our foundation model. Therefore we developed approaches enabling to immediately increase accuracy through emulation while reducing time to full quantum hardware adoption. A follow-up to this discussion, linked to the reduction in qubits requirements can be found in section 4.3.. Lets us now focus on quantum strategies allowing to access fault tolerant architectures to provide near-exact QPE quantum chemistry computations.
@@ 3.2.2 Scalable Algorithms for Fault-Tolerant Architectures: Towards Quantum Phase Estimation
At the opposite end of the spectrum, Quantum Phase Estimation (QPE) provides ground-state energies at Full-CI accuracy with controllable error assuming fault-tolerant hardware. Within a fixed basis, it yields the exact Full-CI solution using a circuit whose size scales with system size, target accuracy, and initial-state quality. Current resource estimates for chemically relevant problems indicate requirements on the order of billions of gates and thousands of logical qubits [78], placing QPE well beyond near-term hardware capabilities, despite its favorable asymptotic scaling and the prospect of at least polynomial quantum advantage. Ongoing efforts therefore focus on reducing these resource demands, a challenge that is twofold: improving the high-level mapping of chemical problems onto quantum operations, and improving the low-level gate decomposition of the fundamental building blocks of QPE. We identify that leading asymptotic formulations of QPE for chemistry rely on first-quantized plane-wave bases [79], which are natural for periodic systems but poorly suited for strongly correlated or highly localized molecular states. To address this, we introduce the use of adaptive molecular grids as a first-quantized basis for QPE, together with an operator transformation yielding a transcorrelated Hamiltonian whose ground states are smoother in real space [77]. Together, these efforts aim to incorporate more chemical structure and physical insight into fault-tolerant quantum algorithms, helping bridge the gap between formal asymptotic advantage and practical chemical applications.
@@
Table 5: Comparison of our First-Quantized Real-Space Framework vs. Standard VQE
Feature First-Quantized Real-Space Framework (FTQC) Standard VQE (NISQ)
Quantization First-quantized, real-space formulation. Typically second-quantized (orbital-based).
Qubit Scaling Logarithmic with spatial resolution: $\mathcal{O}(\eta\log N)$. Linear with the number of spin-orbitals: $\mathcal{O}(N)$.
Primary Algorithm Deterministic: QPE (Hermitian) or QEVE (Transcorrelated). Heuristic: Variational minimization of energy.
Discretization Non-uniform, molecule-adaptive grids (Voronoi partitions). Fixed basis sets (e.g., Gaussian atom-centered functions).
Cusp Handling Transcorrelated (TC) extension eliminates Coulomb singularities. Limited by basis set; requires many functions to approach CBS limit.
Complexity Two-body operators achieve quadratic scaling $\mathcal{O}(N^{2})$. Often involves $\mathcal{O}(N^{4})$ terms in standard representations.
The second aspect of this challenge concerns optimizing the low-level gate decomposition of QPE building blocks, a rapidly evolving area that closely follows advances in hardware and error correction to guide compilation strategies. We introduced an implementation of the multi-controlled NOT gate whose circuit depth scales only polylogarithmically with the number of control qubits [80], yielding an exponential improvement over previous state-of-the-art. To conclude, let us explain how we can illustrate this type of work by implementing building block on today’s hardware.
@@ 3.2.3 Assessing FTQC Building Blocks on NISQ Hardware
To illustrate how NISQ devices can be used to prototype algorithmic primitives designed for fault-tolerant quantum computing (FTQC), we consider a recent method we introduced for preparing quantum states whose amplitudes follow a polynomial of degree $d$. The construction achieves logarithmic circuit depth in the number of qubits $n$, improving on previous approaches that required linear depth. The method combines three ingredients compatible with FTQC architectures: a block-encoding of an affine diagonal operator using a Pauli-basis decomposition with only $n$ terms, a modified linear-combination-of-unitaries (LCU) implementation enabling logarithmic-depth execution, and a generalized quantum eigenvalue transformation (GQET) to promote the affine operator to an arbitrary polynomial. Because polynomial approximations are ubiquitous in scientific computing, this primitive provides a scalable route to quantum state preparation for many quantum algorithms.
A proof-of-principle implementation [81] was carried out on the 56-qubit Quantinuum H2 trapped-ion processor [82], using 14 qubits and more than 500 primitive gates, demonstrating the feasibility of the approach on present-day hardware (Fig. 8). The measured amplitudes follow the expected polynomial profile, confirming the construction experimentally. This illustrates how NISQ hardware can already be used to experimentally validate FTQC-oriented algorithmic primitives at small scale.
@@
These results highlight both the necessity and the opportunity for tight co-design between chemistry-informed algorithm development and hardware-efficient gate-level compilation.

### 19 Mar 2026 v2

@@ 1 Introduction
This here paper is organized to follow a streamlined, end-to-end drug design workflow utilizing both classical and quantum computing platforms. We begin by discussing the preparation of molecular systems, including the water placement problem in quantum computing. We then analyze quantum algorithms for quantum chemistry as providers of high-accuracy training data for quantum-based Neural Network Potentials. A focus on the Hyperion quantum emulator is provided alongside the discussion on quantum hardware. We then focus on the development of these specific Machine Learning models for Molecular Dynamics simulations, introducing the FeNNix-Bio1 foundation model and its underlying Ignis database, and showing the transition from classical to exact quantum algorithms for training data generation.
Finally, we provide an overview of modern enhanced sampling techniques and their projected quantum speedups. Throughout each section, we contrast classical and quantum implementations while providing detailed performance and resource estimations.
@@ 3.2 Quantum Computing: Overview of Recent Quantum Algorithms for Electronic Structure
3.2.1 Near-Term Hybrid Quantum Algorithms: the Search for Optimal Ansätze
The VQE approximates the molecular ground state by minimizing the energy over a truncated subspace of the Full-CI Hilbert space, in close analogy with classical CI methods. A key difference is that VQE constructs this variational space using an ansatz composed of unitary, exponentially parameterized operators, whereas CI relies on linear expansions in excitation operators. In the VQE workflow, a classical optimizer iteratively updates the parameters to minimize the Hamiltonian expectation value, seeking convergence to the lowest-energy state within the variational subspace defined by the ansatz, ideally approximating the true ground state of the system. Quantum advantage with VQE would then mean that, for some practically relevant problems, the variational space spanned by the quantum circuit provides a better heuristic than classical ansätze [65], assuming efficient parameter optimisation. Then, the central challenge in VQE is ansatz design: it must balance expressivity (large variational space), trainability (enabling efficient parameter optimization), and circuit depth, as NISQ devices allow only limited numbers of qubits and gates due to decoherence. So far, practical implementations continue to face numerous challenges, several of which are outlined below. First, commonly used ansätze can be insufficiently expressive for strongly correlated systems, particularly when higher-order excitations such as triples are required to achieve chemical accuracy. Second, the operator pools from which ansätze are constructed are often far larger than algebraically necessary, resulting in substantial computational overhead. Last but not least, VQE relies on a global, non-convex optimization landscape that is often rough and largely flat, due both to barren plateaus arising from high dimensionality and to excited-state contamination intrinsic to energy minimization, leading to significant convergence difficulties in multi-parameter optimization.
@@ 3.2.1 Near-Term Hybrid Quantum Algorithms: the Search for Optimal Ansätze
First, we consider fixed-ansatz approaches, which employ a predefined circuit structure encompassing tunable single-qubit rotation gates (the rotation angles being the parameters to optimise). A conventional example is the tUCCSD ansatz [66], that takes the form
$U_{\mathrm{tUCCSD}}(\boldsymbol{\theta})=\prod_{\mu\in\mathcal{S}}e^{\theta_{\mu}(\tau_{\mu}-\tau_{\mu}^{\dagger})}\prod_{\nu\in\mathcal{D}}e^{\theta_{\nu}(\tau_{\nu}-\tau_{\nu}^{\dagger})},$
where $\tau_{\mu}$ and $\tau_{\nu}$ are single and double excitation cluster operators, and $\tau-\tau^{\dagger}$ ensures unitarity.
To improve expressivity in strongly correlated regimes, we extended this standard tUCCSD by incorporating triple excitations, leading to a tUCCSDT formulation [67] capable of capturing higher-order correlation effects required for chemical accuracy. Second, to leverage hardware-efficient operators, i.e., operators that admit shallower circuit decompositions than the fermionic excitations in tUCC ansätze, the non-iterative disentangled UCC (NI-DUCC) ansatz was introduced [68], built from a highly expressive network ensuring algebraic completeness. Third, an efficient and systematic method was proposed to generate complete operator pools, sets of operators associated with a given Lie algebra from which any unitary can be constructed [69], a formerly technically demanding task as the number of qubit grows. This advance enables the practical deployment of minimal complete pools (MCPs) within NI-DUCC, referred to as NI-DUCC(MCPs).
Table 1: Comparison of representative Variational Quantum Eigensolver (VQE) methods for quantum chemistry. Criteria: (1) Operator Pool, (2) Ansatz Structure, (3) Cost Function, (4) Operator Selection Criterion, (5) Optimisation, (6) Near-term Hardware Compatibility, (7) Target Chemical Regime, (8) Comments.
Method
tUCCSD 1. Fermionic singles/doubles, $\mathcal{O}(N^{4})$.
3. Energy, $\mathcal{O}(N^{4})$ Pauli strings.
5. Global, non-convex.
7. Small systems, moderate correlation.
2. Fixed.
4. Pre-determined.
6. Weak.
8. Chemically inspired, lacks flexibility.
ADAPT 1. Any, up to $\mathcal{O}(N^{4})$.
3. Energy, $\mathcal{O}(N^{4})$ Pauli strings.
5. Global, non-convex.
7. Small systems, moderate/high correlation.
2. Adaptive, operator-wise.
4. On-the-fly energy gradient.
6. Weak.
8. Barren-plateaus resilience.
GGA 1. Any.
3. Energy, $\mathcal{O}(N^{4})$ Pauli strings.
5. Sequential, deterministic update.
7. Small systems, weak correlation
2. Adaptive, operator-wise.
4. On-the-fly energy line-search.
6. Moderate/strong.
8. High trainability, low expressivity.
Overlap-ADAPT 1. Any.
3. Overlap, $\mathcal{O}(1)$ observable.
5. Global, convex.
7. Strong correlation levels.
2. Adaptive, operator-wise.
4. On-the-fly overlap gradient.
6. Strong.
8. Needs accurate target state.
MB-ADAPT 1. Minimal complete pool, $\mathcal{O}(N)$.
3. Energy, $\mathcal{O}(N^{4})$ Pauli strings.
5. Global, non-convex.
7. Large systems, all correlation levels.
2. Adaptive, batch of operators.
4. On-the-fly energy gradients.
6. Intermediate.
8. Reduced gradient costs w.r.t ADAPT.
NI-DUCC(MCPs) 1. Minimal complete pool, $\mathcal{O}(N)$.
3. Energy, $\mathcal{O}(N^{4})$ Pauli strings.
5. Global, non-convex.
7. Large systems, all correlation levels.
2. Fixed, layered.
4. Pre-determined.
6. Intermediate.
8. Very expressive ansatz.
@@ 7. Small systems, weak correlation
In contrast to fixed-structure ansätze, adaptive approaches such as ADAPT-VQE [70, 71] iteratively construct the circuit by selecting and adding operators from a predefined pool based on a given criterion. This enables the generation of compact, problem-tailored circuits. In this context, MCPs provide a possible choice of operator sets. Building on this idea, we showed that selecting and adding operators in batches [72] significantly reduces the computational cost associated with gradient evaluations, leading to the MB-ADAPT-VQE framework [69].
Taken together, NI-DUCC (with MCPs) and MB-ADAPT-VQE yield performance that is arguably superior to prior state-of-the-art across a range of tested molecular systems in relevant regimes.
To address excited-state contamination and the resulting energy plateaus in ADAPT-VQE, we introduced Overlap-ADAPT-VQE [73], which grows the ansatz via variational projection onto an intermediate-quality reference state rather than direct energy minimization. This projection-based strategy yields significantly more compact circuits and accelerates subsequent convergence. More broadly, Overlap-ADAPT-VQE establishes a unified framework for learning and compressing quantum states as ansätze. We exploited this capability for classical data loading by compressing sparse selected-CI wavefunctions into compact quantum representations that serve as high-quality initial states for both VQE and Quantum Phase Estimation [73, 74], two algorithms whose performance is critically dependent on initialization quality.
@@
We conclude by addressing the practical challenges of VQE on current quantum hardware, in particular the optimization difficulties caused by non-convex landscapes. We introduced Greedy Gradient-Free Adaptive-VQE (GGA-VQE) [75], that exploits analytical structure to convert the global variational optimization into an iterative greedy procedure in which each parameter can be determined locally without gradient estimation. This substantially reduces the classical optimization overhead while producing compact, hardware-aligned ansätze tailored to the target Hamiltonian.
@@
Despite significant fluctuations in the QPU-measured energies, the parameters extracted from hardware measurements remain close to the optimal noiseless values for the first iterations of the algorithm (Fig. 6), indicating that the greedy analytical updates are robust to current processor noise levels. At the same time, recent analyses by Dalton et al. [76] show that achieving a genuine quantum advantage with adaptive VQE would require several orders of magnitude reduction in current gate error rates. GGA-VQE can be viewed as a ”minimal” adaptive strategy for hardware implementation because it removes the costly global non-convex optimization step, which is especially sensitive to noise. More generally, current hardware demonstrations of adaptive VQE remain confined, at best, to very small problems and minimal-basis descriptions, still far from the accuracy required for integration into our foundation model. In this context, emulation plays a central role in the development and validation of advanced methods, enabling algorithmic progress beyond current hardware limitations. We now turn to the longer-term perspective of fault-tolerant quantum computing, where Quantum Phase Estimation offers a route to near-exact quantum chemistry calculations.
@@ 3.2.2 Scalable Algorithms for Fault-Tolerant Architectures: Towards Quantum Phase Estimation
At the opposite end of the spectrum, the Quantum Phase Estimation algorithm provides ground-state energies at Full-CI accuracy with controllable error assuming fault-tolerant hardware. Within a fixed basis, it yields the exact Full-CI solution using a circuit whose size scales with system size, target accuracy, and initial-state quality. Current resource estimates for chemically relevant problems indicate requirements on the order of billions of gates and thousands of logical qubits [78], placing QPE well beyond near-term hardware capabilities, despite its favorable asymptotic scaling and the prospect of at least polynomial quantum advantage. Ongoing efforts therefore focus on reducing these resource demands, a challenge that is twofold: improving the high-level mapping of chemical problems onto quantum operations, and improving the low-level gate decomposition of the fundamental building blocks of QPE. We identify that leading asymptotic formulations of QPE for chemistry rely on first-quantized plane-wave bases [79], which are natural for periodic systems but poorly suited for strongly correlated or highly localized molecular states. To address this, we introduce the use of adaptive molecular grids as a first-quantized basis for QPE, together with an operator transformation yielding a transcorrelated Hamiltonian whose ground states are smoother in real space [77]. Together, these efforts aim to incorporate more chemical structure and physical insight into fault-tolerant quantum algorithms, helping bridge the gap between formal asymptotic advantage and practical chemical applications.
@@
Table 2: Comparison between real-space first-quantized and molecular-orbital second-quantized frameworks
Real-space basis Molecular-Orbital basis
Quantization First quantization Second quantization
Qubit count $\mathcal{O}(\eta\log N)$; logarithmic in basis $\mathcal{O}(N)$; no explicit particle dependence
Hamiltonian terms Typically $\mathcal{O}(N^{2})$; diagonal Coulomb term Typically $\mathcal{O}(N^{4})$;
Quantum algorithms QPE, QEVE (FTQC) VQE (NISQ), QPE (FTQC)
Spatial resolution tricks Adaptive grids, Transcorrelated Hamiltonian Adaptive basis sets (SABS), orbital rotations
A second aspect of this challenge concerns the optimization of low-level gate decompositions for key QPE building blocks. In this direction, we developed low-depth implementations of essential primitives, including a multi-controlled NOT gate with polylogarithmic circuit depth in the number of controls [80], representing an exponential improvement over previous constructions. Similarly, we introduced a logarithmic-depth method for quantum state preparation of polynomials [81].
Although these constructions are primarily designed for fault-tolerant quantum computers, current NISQ devices can already serve as valuable testbeds for validating such algorithmic primitives at small scale. We illustrate this approach below with the case of polynomial state preparation.
@@ 3.2.3 Assessing FTQC Building Blocks on NISQ Hardware
As an example, we consider a method we introduced for preparing quantum states with amplitudes following a polynomial of degree $d$. The construction achieves logarithmic circuit depth in the number of qubits $n$, improving over previous approaches with linear depth. It relies on three FTQC-compatible components: a block-encoding of an affine diagonal operator relying on a modified linear-combination-of-unitaries (LCU) scheme enabling logarithmic-depth execution, and a generalized quantum eigenvalue transformation (GQET) to promote the affine operator to arbitrary degree polynomials. As polynomial approximations are ubiquitous in scientific computing, this primitive provides a scalable route to quantum state preparation across a wide range of quantum algorithms.
A proof-of-principle implementation [81] was performed on the 56-qubit Quantinuum H2 trapped-ion processor [82], using 14 qubits and over 500 primitive gates. The measured amplitudes closely match the expected polynomial profile, confirming the validity of the construction. This experiment demonstrates that NISQ hardware can already serve as a testbed for validating FTQC-oriented primitives at small scale.
@@
These results highlight both the necessity and the opportunity for tight co-design between chemistry-informed algorithm development and efficient quantum primitive operations.